# Theory of In-Plane-Magnetic-Field-Dependent Excitonic Spectra in Atomically Thin Semiconductors

### v1

@@ IV.1 Hybridized Excitonic Energies
On the contrary, for very high magnetic-field strengths and small dark-bright splittings, the growth becomes linear, which can be observed in the case for MoSe2 in Fig. 4(a). This regime can be analytically obtained from Eq. 12, if one assumes that the in-plane magnetic field dominates:
@@
In the following, we discuss the impact of the linewidth on the level repulsion. The dependence on the linewidth difference is particularly pronounced for MoSe2 in Fig. 4(a), because it exhibits a small dark-bright splitting. As indicated by the color gradient, the spin-hybridization at a fixed magnetic field strength varies, if the linewidth difference is varied: If $\kappa$ increases, the degree of hybridization and hence the splitting decreases. Therefore, the dissipation of the system effectively suppresses spin-hybridization and can also be viewed as an effective attractive interaction.
@@
In this case, the initial quadratic growth becomes exclusively inversely proportional to the square of $\kappa$. This holds true for MoSe2 in Fig. 4(a). (ii) On the other hand, when the dark-bright splitting $\Delta$ becomes the dominating parameter, which holds true for the majority of the TMDCs (MoS2, WS2, WSe2), we obtain an expression that does not depend on $\kappa$ anymore:
@@
This case, for both materials in Figs. 4(a) and 4(b), leads to the strongest shifts, as the attractive part of the coupling induced by the dissipation is absent. Eq. 29 is the usual expression of level repulsion [55] and has been already discussed in other works [19, 18]. We note, that the magnetic-field-induced hybridization also manifests in time domain via quantum beats [56].
@@ IV.2 Hybridized Excitonic Linewidths
The hybridized linewidths $\hbar\gamma_{B_{\parallel}}^{\mathcal{S}}$ are governed by Eq. 13 and plotted for MoSe2 in Fig. 5. We note, that the linewidths discussed in this manuscript always correspond to a coherence decay via, e.g., radiative or pure dephasing and are not equal to a population decay due to radiative or nonradiative recombination [57].
@@ IV.3 Absorption ratios of hybridized exciton peaks
In Fig. 7(a) and Fig. 7(b), we show the peak amplitude ratios from Eq. 34 for the materials MoS2 and MoSe2, respectively. We evaluate Eq. 34 in such a way, that the numerator (denominator) always corresponds to the spin-dark (spin-bright) resonance, and we choose a regime of quasi radiatively-limited linewidth, which corresponds to disorder-/defect-free h-BN-encapsulated samples at cryogenic temperatures.
@@
Figure 7: Amplitude peak ratios and spin-bright/spin-dark amplitudes, that are normalized w.r.t. $\mathcal{N}=\alpha(\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}})+\alpha(\hbar\omega_{x,B_{\parallel}}^{\overline{\mathcal{S}}})$, for MoS2 (Fig. 7(a)) and MoSe2 (Fig. 7(b)) with non-radiative linewidth $\hbar\gamma_{\text{nrad}}=0.1$ meV for increasing in-plane magnetic field-strengths. The colourgradient of the curves show the values of the real parts of mixing coefficients from Eq. 18, i.e. the hybridization coefficients. Red (blue) represents a spin-up (spin-down) electron-state and yellow denotes the degree of spin-mixing.
@@
This approximate expression, cf. silver dashed line in Fig. 7(a), describes the initial grow with respect to the magnetic-field strength, which reproduces the full expression, cf. olive solid line, very well. For magnetic fields above 30 T, however, LABEL:eq:_peak_ratio_MoS2 fails to reproduce the correct saturation behavior: It asymptotically approaches a ratio exceeding 50 $\%$, whereas the exact expression exhibits a maximum before converging to 50 $\%$ as illustrated for MoSe2 in Fig. 7(b). We further plot the quadratic approximation of LABEL:eq:_peak_ratio_MoS2, cf. silver dotted line in Fig. 7(a), with respect to the excitonic magnetic matrix element $\mathcal{B}^{2}$ around the origin, which reproduces the full expression (olive solid line) up to roughly 15 T. We note that the limit case in LABEL:eq:_peak_ratio_MoS2 can be used to model any material that exhibits a dominating dark-bright splitting over the linewidths, for example WS2 and WSe2 at cryogenic temperatures.
In contrast, within the field strengths considered here, the peak amplitude ratio for a MoSe2 monolayer, cf. olive solid line in Fig. 7(b), displays a strong non-monotonous behavior at increasing magnetic-field strengths. Initially, it increases, until a turning point at around 13 T is reached, from where it then slowly decreases. This non-monotonous behavior is a direct consequence of the fact that in a MoSe2 monolayer, dark-bright splitting, linewidths and linewidths differences are all of similar magnitude in the range of 1–2 meV. The strong initial growth of the peak ratio well beyond 100 $\%$ can be explained by the fact that, due to the absence of a radiative contribution, the linewidth of the spin-dark resonance is smaller compared to the spin-bright resonance. This in turn overemphasizes the oscillator strength it receives from hybridization (smaller linewidths lead to larger peak amplitudes at a constant area under the curve). For a closer analysis, we rewrite the peak ratio from Eq. 34 as:
@@
which is a product of a monotonously decreasing ratio of the hybridized linewidths $\frac{(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}=-1})^{2}}{(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}=1})^{2}}$ and a monotonously increasing function $g^{\mathcal{S}}$, defined in Eq. 42, with respect to the magnetic field $B_{\parallel}$. Hence, the maximum of the amplitude ratio in Fig. 7(b) can be rationalized as that point, where the linewidth ratio $\frac{(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}=-1})^{2}}{(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}=1})^{2}}$ reflecting the magnetic-field-induced linewidth redistribution starts to grow faster than the oscillator-strength redistribution encoded in $g^{\mathcal{S}}$. At magnetic fields exceeding the turning point, a decrease of the amplitude ratio occurs, until it is fully determined by the degree of the spin hybridization, as $\frac{(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}=-1})^{2}}{(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}=1})^{2}}\overset{B_{\parallel}\rightarrow\infty}{\longrightarrow}1$ and $g^{\mathcal{S}}\overset{B_{\parallel}\rightarrow\infty}{\longrightarrow}0.5$. Consequently, if the linewidths are initially equal, the linewidth ratio becomes unity and the the peak ratio in Eq. 36 will be a purely monotonic increase with no extremum, as in the case for MoS2 in LABEL:eq:_peak_ratio_MoS2 (within the field strengths considered).
Note, that within our definition, cf. Eq. 34, the spin-dark amplitude and, thus, the ratio are not exactly zero for a vanishing in-plane magnetic field, cf. Fig. 7(b), because the spin-dark resonance lies in close vicinity to the spin-bright resonance.
We point out, that the qualitative behavior of the peak ratios resembles the behavior of the peak ratios extracted from PL experiments [20, 19]. However, while similar in their outer appearance, the underlying physics is different: In contrast to absorption measurements, which display the coherent response of the material, PL displays the emission of radiation. Hence, the PL amplitudes always scale with the incoherent excitonic occupations at the respective exciton resonance, which recombine within the light cone [57, 58, 52, 32], and the peak ratio can be rationalized as the ratio of the Boltzmann-distributed spin-bright and spin-dark occupations [19]. Therefore, future work can possibly shed more light on this issue, i.e., on the question, if the non-monotonous behavior can also be observed in actual absorption measurements.
@@ Conclusion
We derived a fully analytical model for the absorption spectrum of the excitonic spin-dark and spin-bright ground states under spin hybridization of an applied in-plane magnetic field. We found, that an intricate interplay of hybridized excitonic energies, linewidths and amplitudes governs the linear optical response. In particular, we showed that a MoSe2 monolayer with quasi radiatively-limited linewidths gives rise to a non-monotonous amplitude behavior of spin-bright and spin-dark excitonic states in optical absorption spectra, similar to recent photoluminescence experiments. The derived equations Eqs. 24, 12 and 13, can be used to analyze magneto-optical experiments in the field of TMDC excitons.
@@ Acknowledgements.
The authors acknowledge financial support by the Deutsche Forschungsgemeinschaft (DFG) through Project No. 420760124.
Appendix A Analytical Formulas
@@
We plug in the definition from Eq. 37 and reorganize the expression using partial fraction decomposition, which leads to the absorption spectrum given in Eq. 24. The amplitudes are given by:
$\mathcal{F}^{\mathcal{S}}$ $=-2\Big[|(i\hbar\gamma_{\text{rad}}^{b})P^{\mathcal{S}}_{B_{\parallel}}|^{2}$ (39)
$\qquad+\Re[(\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}}+i\hbar\gamma_{B_{\parallel}}^{\mathcal{S}})(P^{\mathcal{S}}_{B_{\parallel}}i\hbar\gamma_{\text{rad}}^{b}+2\mathcal{S}\Lambda)]\Big],$
$\mathcal{G}^{\mathcal{S}}$ $=-2\Re\Big[i\hbar\gamma_{\text{rad}}^{b}P^{\mathcal{S}}_{B_{\parallel}}+2\mathcal{S}\Lambda\Big],$ (40)
@@
$\Lambda=\frac{(\hbar\gamma_{\text{rad}}^{b})^{2}P^{\mathcal{S}=1}_{B_{\parallel}}P^{*\,\mathcal{S}=-1}_{B_{\parallel}}}{\hbar\omega_{x,B_{\parallel}}^{{\mathcal{S}=-1}}+i\hbar\gamma_{B_{\parallel}}^{{\mathcal{S}=-1}}-\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}=1}+i\hbar\gamma_{B_{\parallel}}^{\mathcal{S}=1}}.$ (41)
The function $g^{\mathcal{S}}$ from Eq. 36 reads:
$\begin{split}g^{\mathcal{S}}=&\frac{\big((\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}}-\hbar\omega_{x,B_{\parallel}}^{\overline{\mathcal{S}}})^{2}+(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}})^{2}\big)}{\big((\hbar\omega_{x,B_{\parallel}}^{\overline{\mathcal{S}}}-\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}})^{2}+(\hbar\gamma_{B_{\parallel}}^{\overline{\mathcal{S}}})^{2}\big)}\\ &\times\frac{(\mathcal{F}^{\mathcal{S}}-\mathcal{G}^{\mathcal{S}}\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}})\big((\hbar\omega_{x,B_{\parallel}}^{\overline{\mathcal{S}}}-\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}})^{2}+(\hbar\gamma_{B_{\parallel}}^{\overline{\mathcal{S}}})^{2}\big)+(\mathcal{F}^{\overline{\mathcal{S}}}-\mathcal{G}^{\overline{\mathcal{S}}}\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}})(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}})^{2}}{(\mathcal{F}^{\mathcal{S}}-\mathcal{G}^{\mathcal{S}}\hbar\omega_{x,B_{\parallel}}^{\overline{\mathcal{S}}})(\hbar\gamma_{B_{\parallel}}^{\overline{\mathcal{S}}})^{2}+(\mathcal{F}^{\overline{\mathcal{S}}}-\mathcal{G}^{\overline{\mathcal{S}}}\hbar\omega_{x,B_{\parallel}}^{\overline{\mathcal{S}}})\big((\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}}-\hbar\omega_{x,B_{\parallel}}^{\overline{\mathcal{S}}})^{2}+(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}})^{2}\big)}.\end{split}$ (42)

### v2

@@ IV.1 Hybridized Excitonic Energies
On the contrary, for very high magnetic-field strengths and small dark-bright splittings, the growth becomes linear, which can be observed in the case for MoSe2 in olive Fig. 4(a). This regime can be analytically obtained from Eq. 12, if one assumes that the in-plane magnetic field dominates:
@@
In the following, we discuss the impact of the linewidth difference on the level repulsion. The dependence on the linewidth difference is particularly pronounced for MoSe2 in Fig. 4(a), because it exhibits a small dark-bright splitting. As indicated by the color gradient, the spin-hybridization at a fixed magnetic field strength varies, if the linewidth difference is varied: If $\kappa$ increases, the degree of hybridization and hence the splitting decreases. Therefore, the dissipation of the system effectively suppresses spin-hybridization and can also be viewed as an effective attractive interaction.
@@
In this case, the initial quadratic growth becomes exclusively inversely proportional to the square of $\kappa$. This holds true for MoSe2 in Fig. 4(a) for dominating $\kappa$. (ii) On the other hand, when the dark-bright splitting $\Delta$ becomes the dominating parameter, which holds true for the majority of the TMDCs (MoS2, WS2, WSe2), we obtain an expression that does not depend on $\kappa$ anymore:
@@
This case, for both materials in Fig. 4(a)–(b), leads to the strongest shifts, as the attractive part of the coupling induced by the dissipation is absent. Eq. 30 is the usual expression of level repulsion [51] and has been already discussed in other works [39, 18]. In Tab. 1, we summarize the measured level repulsion at 15 T and 30 T and compare these values with our model according to Eq. 12; we observe a good agreement between experiment and theory. We suspect that small differences at 30 T are due to varying linewidth differences $\kappa$ for both experiments.
Ref. Material $B_{\parallel}=15$ T (Exp./Th.) $B_{\parallel}=30$ T (Exp./Th.)
[50] MoSe2 0.4 / 0.4 meV 1.2 / 1.06 meV
[39] MoSe2 0.4 / 0.4 meV 1.1 / 1.06 meV
Table 1: Level repulsion of experiments (Exp.) and our theory at $\kappa=0$ (Th.) at 15 T and 30 T.
We note that the magnetic-field-induced hybridization also manifests in time domain via quantum beats [41].
@@ IV.2 Hybridized Excitonic Linewidths
The hybridized linewidths $\hbar\gamma_{B_{\parallel}}^{\mathcal{S}}$ are governed by Eq. 13 and plotted for MoSe2 in Fig. 5. We note that the linewidths, discussed in this manuscript, always correspond to a coherence decay via, e.g., radiative or pure dephasing and are not equal to a population decay due to radiative or nonradiative recombination [56].
@@ IV.3 Absorption ratios of hybridized exciton peaks
In Fig. 7(a) and Fig. 7(b), we show the peak amplitude ratios from Eq. 35 for the materials MoS2 and MoSe2, respectively. We evaluate Eq. 35 in such a way that the numerator (denominator) always corresponds to the spin-dark (spin-bright) resonance, and we choose a regime where the radiative and non-radiative linewidths are roughly on the same scale, which corresponds to disorder-/defect-free h-BN-encapsulated samples at cryogenic temperatures.
@@
Figure 7: Spin-bright/spin-dark amplitudes, that are normalized w.r.t. $\mathcal{N}=\alpha(\hbar\omega_{x,B_{\parallel}=0}^{\mathcal{S}})+\alpha(\hbar\omega_{x,B_{\parallel}=0}^{\overline{\mathcal{S}}})$ and amplitude peak ratios for MoS2 ( Fig. 7(a)) and MoSe2 (Fig. 7(b)) with non-radiative linewidth $\hbar\gamma_{\text{nrad}}=0.1$ meV for increasing in-plane magnetic field-strengths. The colourgradient of the curves show the values of the real parts of mixing coefficients from Eq. 18, i.e. the hybridization coefficients. Red (blue) represents a spin-up (spin-down) electron-state and yellow denotes the degree of spin-mixing.
@@
This approximate expression, cf. silver dashed line in Fig. 7(a), describes the initial grow with respect to the magnetic-field strength, which reproduces the full expression, cf. olive solid line, very well. For magnetic fields above 30 T, however, LABEL:eq:_peak_ratio_MoS2 fails to reproduce the correct saturation behavior: It asymptotically approaches a ratio exceeding 100 $\%$, whereas the exact expression exhibits a maximum before converging to 100 $\%$ as illustrated for MoSe2 in Fig. 7(b). We further plot the quadratic approximation of LABEL:eq:_peak_ratio_MoS2, cf. silver dotted line in Fig. 7(a), with respect to the excitonic magnetic matrix element $\mathcal{B}^{2}$ around the origin, which reproduces the full expression (olive solid line) up to roughly 15 T. We note that the limit case in LABEL:eq:_peak_ratio_MoS2 can be used to model any material that exhibits a dominating dark-bright splitting over the linewidths, for example WS2 and WSe2 at cryogenic temperatures.
In contrast, within the field strengths considered here, the peak amplitude ratio for a MoSe2 monolayer, cf. olive solid line in Fig. 7(b), displays a strong non-monotonous behavior at increasing magnetic-field strengths. Initially, it increases, until a turning point at around 10 T is reached, from where it then slowly decreases. This non-monotonous behavior is a direct consequence of the fact that in a MoSe2 monolayer, dark-bright splitting, linewidths and linewidths differences are all of similar magnitude in the range of 1–2 meV. The strong initial growth of the peak ratio well beyond 100 $\%$ can be explained by the fact that, due to the absence of a radiative contribution, the linewidth of the spin-dark resonance is smaller compared to the spin-bright resonance. This in turn overemphasizes the oscillator strength it receives from hybridization (smaller linewidths lead to larger peak amplitudes at a constant area under the curve). For a closer analysis, we rewrite the peak ratio from Eq. 35 as:
@@
which is a product of a monotonously decreasing ratio of the hybridized linewidths $\frac{(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}=-1})^{2}}{(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}=1})^{2}}$ and a monotonously increasing function $g^{\mathcal{S}}$ with respect to the magnetic field $B_{\parallel}$, defined as
$\begin{split}&g^{\mathcal{S}}=\frac{\big[(\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}}-\hbar\omega_{x,B_{\parallel}}^{\overline{\mathcal{S}}})^{2}+(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}})^{2}\big]}{\big[(\hbar\omega_{x,B_{\parallel}}^{\overline{\mathcal{S}}}-\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}})^{2}+(\hbar\gamma_{B_{\parallel}}^{\overline{\mathcal{S}}})^{2}\big]}\\ &\times\frac{\mathcal{A}^{\mathcal{S}}_{B_{\parallel}}\big[(\hbar\omega_{x,B_{\parallel}}^{\overline{\mathcal{S}}}-\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}})^{2}+(\hbar\gamma_{B_{\parallel}}^{\overline{\mathcal{S}}})^{2}\big]+\mathcal{A}^{\overline{\mathcal{S}}}_{B_{\parallel}}(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}})^{2}}{\mathcal{A}^{\overline{\mathcal{S}}}_{B_{\parallel}}\big[(\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}}-\hbar\omega_{x,B_{\parallel}}^{\overline{\mathcal{S}}}]^{2}+(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}})^{2}\big]+\mathcal{A}^{\mathcal{S}}_{B_{\parallel}}(\hbar\gamma_{B_{\parallel}}^{\overline{\mathcal{S}}})^{2}}.\end{split}$ (38)
Hence, the maximum of the amplitude ratio in Fig. 7(b) can be rationalized as that point, where the linewidth ratio $\frac{(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}=-1})^{2}}{(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}=1})^{2}}$ reflecting the magnetic-field-induced linewidth redistribution starts to grow faster than the oscillator-strength redistribution encoded in $g^{\mathcal{S}}$. At magnetic fields exceeding the turning point, a decrease of the amplitude ratio occurs, until it is fully determined by the degree of the spin hybridization, as $\frac{(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}=-1})^{2}}{(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}=1})^{2}}\overset{B_{\parallel}\rightarrow\infty}{\longrightarrow}1$ and $g^{\mathcal{S}}\overset{B_{\parallel}\rightarrow\infty}{\longrightarrow}$ 1. Consequently, if the linewidths are initially equal, the linewidth ratio becomes unity and the the peak ratio in Eq. 37 will be a purely monotonic increase with no extremum, as in the case for MoS2 in LABEL:eq:_peak_ratio_MoS2 (within the field strengths considered). Note that within our definition, cf. Eq. 35, the spin-dark amplitude and, thus, the ratio are not exactly zero for a vanishing in-plane magnetic field, cf.  Fig. 7(b), because the spin-dark resonance lies in close vicinity to the spin-bright resonance.
We point out that the qualitative behavior of the peak ratios resembles the behavior of the peak ratios extracted from PL experiments [50, 39]. However, while similar in their outer appearance, the underlying physics is different: In contrast to absorption measurements, which display the coherent response of the material, PL displays the emission of radiation. Hence, the PL amplitudes always scale with the incoherent excitonic occupations at the respective exciton resonance, which recombine within the light cone [56, 4, 19, 45], and the peak ratio can be rationalized as the ratio of the Boltzmann-distributed spin-bright and spin-dark occupations [39]. Therefore, future work can possibly shed more light on this issue, i.e., on the question, if the non-monotonous behavior can also be observed in actual absorption measurements.
@@ Conclusion
We derived a fully analytical model for the absorption spectrum of the excitonic spin-dark and spin-bright ground states under spin hybridization of an applied in-plane magnetic field. We found, that an intricate interplay of hybridized excitonic energies, linewidths and amplitudes governs the linear optical response. In particular, we showed that a MoSe2 monolayer with sufficiently small linewidths gives rise to a non-monotonous amplitude behavior of spin-bright and spin-dark excitonic states in optical absorption spectra, similar to recent photoluminescence experiments. The derived equations Eqs. 24, 12 and 13, can be used to analyze magneto-optical experiments in the field of TMDC excitons.
Data availability statement
All data that support the findings of this study are included within the article.
@@ Acknowledgements.
The authors acknowledge financial support by the Deutsche Forschungsgemeinschaft (DFG) through Project No. 420760124. H.M. acknowledges funding by project 21209528 (“proof of trust”).
Appendix A Absorption spectrum
@@
We plug in the definition from Eq. 39 and reorganize the expression using partial fraction decomposition, which leads to the following absorption spectrum
$\alpha(\omega)=\sum_{\mathcal{S}}\frac{\mathcal{F}^{\mathcal{S}}_{B_{\parallel}}+\mathcal{G}^{\mathcal{S}}_{B_{\parallel}}(\hbar\omega-\hbar\omega_{x}^{\mathcal{S}})}{(\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}}-\hbar\omega)^{2}+(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}})^{2}},$ (41)
The amplitudes are given by:
$\mathcal{F}^{\mathcal{S}}_{B_{\parallel}}$ $=-2(\hbar\gamma_{\text{rad}}^{b})^{2}|P^{\mathcal{S}}_{B_{\parallel}}|^{2}$ (42)
$\qquad+2\hbar\gamma_{B_{\parallel}}^{\mathcal{S}}\big(\hbar\gamma_{\text{rad}}^{b}\text{Re}[P^{\mathcal{S}}_{B_{\parallel}}]+2\mathcal{S}\text{Im}[\Lambda_{B_{\parallel}}]\big),$
$\mathcal{G}^{\mathcal{S}}_{B_{\parallel}}$ $=2\Re\Big[i\hbar\gamma_{\text{rad}}^{b}P^{\mathcal{S}}_{B_{\parallel}}+2\mathcal{S}\Lambda_{B_{\parallel}}\Big],$ (43)
@@
$\Lambda_{B_{\parallel}}=\frac{(\hbar\gamma_{\text{rad}}^{b})^{2}P^{\mathcal{S}=1}_{B_{\parallel}}P^{*\,\mathcal{S}=-1}_{B_{\parallel}}}{\hbar\omega_{x,B_{\parallel}}^{{\mathcal{S}=-1}}-\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}=1}+i(\hbar\gamma^{b}+\hbar\gamma^{d})}.$ (44)
To obtain a simplified formula for the linear absorption spectrum, we can make the following close-to-resonance approximation
$\mathcal{G}^{\mathcal{S}}_{B_{\parallel}}\cdot(\hbar\omega-\hbar\omega_{x}^{\mathcal{S}})\approx 0$ (45)
yielding the linear absorption spectrum
$\alpha(\omega)=\sum_{\mathcal{S}}\frac{\mathcal{A}^{\mathcal{S}}_{B_{\parallel}}}{(\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}}-\hbar\omega)^{2}+(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}})^{2}},$ (46)
which is a superposition of a pair of Lorentzians.